# JAX-SPH: A Differentiable Smoothed Particle Hydrodynamics Framework

### v1

@@ Abstract
Particle-based fluid simulations have emerged as a powerful tool for solving the Navier-Stokes equations, especially in cases that include intricate physics and free surfaces. The recent addition of machine learning methods to the toolbox for solving such problems is pushing the boundary of the quality vs. speed tradeoff of such numerical simulations. In this work, we lead the way to Lagrangian fluid simulators compatible with deep learning frameworks, and propose JAX-SPH – a Smoothed Particle Hydrodynamics (SPH) framework implemented in JAX. JAX-SPH builds on the code for dataset generation from the LagrangeBench project (Toshev et al., 2023b) and extends this code in multiple ways: (a) integration of further key SPH algorithms, (b) restructuring the code toward a Python library, (c) verification of the gradients through the solver, and (d) demonstration of the utility of the gradients for solving inverse problems as well as a Solver-in-the-Loop application. Our code is available at https://github.com/tumaer/jax-sph.
@@ 1 Introduction
Partial differential equations (PDEs) are the mathematical tools developed to describe natural phenomena ranging from engineering and physics to social sciences. Various numerical methods have been developed to solve these PDEs, as analytical solutions are only available for toy examples, with the most recent class of PDE solvers being the machine learning-based ones (Thuerey et al., 2021; Brunton & Kutz, 2023). One particular class of machine learning (ML) approaches called hybrid solvers refers to combining ideas (or full algorithmic blocks) from classical numerical solvers and machine learning (Schenck & Fox, 2018; Um et al., 2020; Kochkov et al., 2021; Jagtap et al., 2022; Lienen & Günnemann, 2022; Karlbauer et al., 2022; Li & Farimani, 2022; Toshev et al., 2023c).
This has been one of the main reasons for the development of differentiable fluid mechanics solvers like PhiFlow (Holl et al., 2020), JAX-CFD (Kochkov et al., 2021), and JAX-Fluids (Bezgin et al., 2022). However, these three frameworks implement Eulerian, i.e., grid-based, solvers, and we lead the way to a JAX-based Lagrangian CFD solver. Eulerian solvers refer to numerical methods that discretize space into static volume elements and then track the evolution of the fluid properties at these positions, while Lagrangian solvers discretize individual material elements, which are then shifted in space following the local velocity field.
@@
As SPH techniques advance, various software packages have arisen in the past few years. However, most of them are designed for high-performance computing (HPC) applications and are typically implemented in low-level languages like C++ or Fortran (Crespo et al., 2015; Koschier et al., 2019; Zhang et al., 2021), with two notable exceptions: PySPH (Ramachandran et al., 2021) in Python and juSPH (Luo et al., 2022) in Julia. Nevertheless, hardly an SPH implementation exists which is out-of-the-box compatible with modern deep learning frameworks like TensorFlow (Abadi et al., 2015), PyTorch [^1] For similar work in progress, we refer the reader to TorchSPH https://github.com/wi-re/pytorchSPH. Compared to our JAX-SPH, this repo uses PyTorch, different SPH algorithms, different experiments, and does not validate the solver or the gradients. (Paszke et al., 2019) or JAX (Bradbury et al., 2018), i.e., leverages automatic differentiation to power differentiable solvers like PhiFlow (Holl et al., 2020). Upon identifying the lack of an ML-ready SPH solver, we choose JAX for the framework implementation for two reasons: (a) JAX tends to be faster for operations on graphs even after PyTorch 2.0 has been introduced (see appendix F in Toshev et al. (2023b)), and (b) we can use the cell list-based neighbor search routine of the JAX-MD library (Schoenholz & Cubuk, 2020). We note that with our solver, we target the easier integration of SPH with ML workflows rather than developing a parallel HPC code, and we exclusively use the AD routine grad by JAX. We leave the implementation of better custom adjoints along the lines of Ma et al. (2021); Kidger (2022); Nadarajah & Jameson (2000) to future work.
In this work, we significantly extend the codebase used for dataset generation within LagrangeBench [^2] https://github.com/tumaer/lagrangebench/blob/main/notebooks/data_gen.ipynb. (Toshev et al., 2023b) and demonstrate the utility of the obtained gradients in multiple ways. Our contributions are:
@@
  • Validating the accuracy of the automatic differentiation-based gradients (Griewank & Walther, 2008) over a few solver steps.
  • A restructured code package, which we will open-source as a Python library.
@@ 2 SPH Solver
In this section, we introduce the components included in our solver code, and in appendix A, we include various solver validation cases.
@@
The preferred way of estimating the density (Monaghan, 2005) is through density summation $\rho_{i}=\sum_{j}m_{j}W(r_{ij}|h)$, with $m_{j}$ being the mass of the adjacent particles $j$, $h$ the smoothing length of the kernel, and $r_{ij}$ the interparticle distance. However, this approach leads to unphysically low densities at free surfaces, forcing the use of density evolution, i.e., numerically integrating the mass conservation Eq. 1. We also implement a density field reinitialization method  (Zhang et al., 2017a) to mitigate errors arising from density evolution (Colagrossi & Landrini, 2003). In weakly compressible SPH, the pressure is defined as a function of density through an equation of state, and in our code, we choose the formulation by Monaghan (1994).
@@
Riemann SPH. We also add the Riemann SPH solver formulation by Zhang et al. (2017b), which is based on the idea of introducing a simple low dissipation limiter to a classical Riemann solver, resulting in decreased numerical dissipation. We construct an inter-particle Riemann problem along a unit vector pointing from particle $i$ to particle $j$, which implicitly regularizes both the momentum equation and the mass conservation. We implement the dissipation limiter as proposed by Zhang et al. (2017b).
Wall boundaries. We follow the generalized wall boundary condition approach by Adami et al. (2012) to enforce different boundary conditions. This approach, in essence, implements two physical constraints: (a) impermeability of walls and (b) viscous effects at walls. For impermeability, the pressure of the adjacent fluid particles is assigned to the dummy wall particles, which leads to a zero pressure gradient in the wall-normal direction at the wall surface, and, thus prevents the penetration of fluid particles into the wall. Regarding viscosity, there are two cases to distinguish: (b1) no-slip enforces the no-slip boundary condition, i.e., the fluid must have zero velocity directly at the wall surface, and (b2) free-slip, i.e., the fluid must have a zero velocity normal to a wall, but might have any velocity tangentially. These two conditions are implemented by assigning the inverted velocity of the fluid to the wall particles, either fully for (b1) or only in the wall-normal direction for (b2). See Adami et al. (2012) for more details.
@@ 3 Experiments
Gradient validation. To test the validity of the gradients obtained through our differentiable solver, we compute analytical solver gradients via automatic differentiation and compare those to numerical gradients from finite difference schemes (Griewank & Walther, 2008). In our setup, gradients are accumulated over 5 solver steps, preceded by 10 (forward only) state warm-up steps. Epsilon for finite differences is picked as $50.001dx=5e-$ 5, as smaller values lead to instabilities. Fig. 1 shows the scalar gradients of kinetic energy over position changes $\frac{dE_{kin}}{dr}$ when using 2-dimensional Taylor-Green vortex (TGV) (Brachet et al., 1983; 1984) and lid-driven cavity (LDC) (Ghia et al., 1982) as initial states.
@@
Inverse problem. Our first application case is an inverse problem, representing the class of inverse design and flow control problems that are successfully tackled by differentiable solvers (Holl et al., 2020) as well as differentiable learned solver surrogates (Allen et al., 2022). The scenario involves a 2D box containing a water cube, discretized by 36 particles, which undergoes acceleration due to gravity across 100 solver steps. The task is to find the initial coordinates of the cube if the final coordinates are given; see Fig. 2.
@@ 4 Conclusion
We have developed JAX-SPH, a framework for simulating Lagrangian fluid problems, which can be easily integrated into design/control problems as well as hybrid solver approaches. By building our code on the high-performance library JAX and validating the simulation results of our solver, we offer a fast and reliable SPH solver in Python. With our work, we hope to accelerate the development of more hybrid Lagrangian solvers, e.g.,  Toshev et al. (2024), and we leave the addition of more SPH algorithms and simulation cases to future work. One particularly exciting future direction is developing foundation models for PDEs that can operate on both Eulerian and Lagrangian data (Alkin et al., 2024), potentially combined with encoded symmetries (Toshev et al., 2023a).
@@ A.1 Taylor Green Vortex 2D
Fig. 4 shows the absolute velocities of each particle at the start and at the end of the Taylor Green vortex (Brachet et al., 1983; 1984) simulation at $Re=100$. Here, the transport velocity formulation of SPH is used. In Fig. 5, one can see the decay of the maximum velocity and the kinetic energy over time for the different solvers, standard SPH (SPH), transport velocity formulation SPH (SPH + tvf), and Riemann SPH (Riemann).
Figure 4: 2D Taylor Green vortex velocity magnitudes at the start of the simulation $t=0$ (left) and at $t=5$ (right), calculated using transport velocity formulation SPH.
Figure 5: 2D Taylor Green vortex solver comparison for $u_{max}$ (left) and $E_{kin}$ (right) between standard SPH, transport velocity formulation SPH, and Riemann SPH at $dx=0.02$ and $dx=0.01$.
@@ A.2 Lid-Driven Cavity 2D
The following figures compare a 2D lid-driven cavity (Ghia et al., 1982) simulation for the different solvers at $Re=100$. The reference data for the velocity profiles, i.e., black dots for U (velocity in $x$ direction at a vertical cut through the middle of the cavity) and black squares for V (velocity in $y$ direction at a horizontal cut through the middle of the cavity), are from Ghia et al. (1982).
Figure 6: Lid-driven cavity with $dx=0.02$ showing absolute particle velocities of the Riemann solver (left) and velocity profiles of each solver at the midsection for U and V (right)
Figure 7: Lid-driven cavity with $dx=0.01$ showing absolute particle velocities of the Riemann solver (left) and velocity profiles of each solver at the midsection for U and V (right)
@@ A.3 Dam Break 2D
The following Fig. 8 shows the nondimensionalized pressure of a Riemann SPH (Zhang et al., 2017b) dam break (Colagrossi & Landrini, 2003) simulation at different time stamps. The fluid flows from the initial state on the left to the right side, interacts with the wall, and reflects a wave backward throughout the domain. A similar validation plot for the standard SPH formulation of our solver is presented in Toshev et al. (2023b).
@@ A.4 Taylor-Green Vortex 3D
Fig. 9 compares the solvers on 3D TGV, similarly to Fig. 5 for 2D. Again, the Reynolds number is set to $Re=100$, and the number of particles in the unit cube is $20^{3}$, $32^{3}$, and $50^{3}$, leading to $dx=0.314$, $dx=0.196$, and $dx=0.126$, respectively. The reference solution is obtained using JAX-Fluids Bezgin et al. (2022) with a $128^{3}$ grid.
Figure 9: 3D Taylor Green vortex solver comparison for $dE/dt$ (left) and $E_{kin}$ (right) between standard SPH, transport velocity formulation SPH, and Riemann SPH with $dx=0.314$, $dx=0.196$, and $dx=0.126$
@@ B.1 Implementation details
Given the number of temporal coarsening steps $M$ and the number of SitL calls n_sitl, each call counts for $\frac{M}{\texttt{n\_sitl}}$ actual timesteps. This results in two distinct time steps ($dt$), necessitating the transformation of input properties to the SitL physical space.
1def sitl_forward(self, r, u_M_norm, nbrs_GNN, dt, M, n_sitl, sigma_uM, sigma_aM):
@@ 9        M: level of temporal coarsening
15    u_phys = u_M_norm * sigma_uM / (dt * M)  # to physical
16    nbrs_SPH = self.nbrs_SPH_update(r)  # internal neighbor list
@@
34        nbrs_GNN = nbrs_GNN_update(r_new)
@@
39    # finite differences to get M-step effective quantities
@@ B.2 Training and results
Solver-in-the-Loop was trained with LagrangeBench (Toshev et al., 2023b), with a custom RPF 2D dataset with 20-step temporal coarsening. For SitL, n_sitl=3. Both the corrector model and the reference GNS (Sanchez-Gonzalez et al., 2020) are message-passing networks with 10 layers and 64 latent dimensions. The starting learning rate is set to $31e-$ 3 for both, and noise_std is set to $51e-$ 5 for SitL and to $31e-$ 3 for GNS. Table 1 shows the LagrangeBench performance metrics on these models. Best models are picked based on the $\text{MSE}_{20}$ loss on the validation set.

### v2

@@ Abstract
Particle-based fluid simulations have emerged as a powerful tool for solving the Navier-Stokes equations, especially in cases that include intricate physics and free surfaces. The recent addition of machine learning methods to the toolbox for solving such problems is pushing the boundary of the quality vs. speed tradeoff of such numerical simulations. In this work, we lead the way to Lagrangian fluid simulators compatible with deep learning frameworks, and propose JAX-SPH – a Smoothed Particle Hydrodynamics (SPH) framework implemented in JAX. JAX-SPH builds on the code for dataset generation from the LagrangeBench project (Toshev et al., 2024b) and extends this code in multiple ways: (a) integration of further key SPH algorithms, (b) restructuring the code toward a Python package, (c) verification of the gradients through the solver, and (d) demonstration of the utility of the gradients for solving inverse problems as well as a Solver-in-the-Loop application. Our code is available at https://github.com/tumaer/jax-sph.
@@ 1 Introduction
Partial differential equations (PDEs) are the mathematical tools developed to describe natural phenomena ranging from engineering and physics to social sciences. Various numerical methods have been developed to solve these PDEs, as analytical solutions are only available for toy examples, with the most recent class of PDE solvers being the machine learning-based ones (Thuerey et al., 2021; Brunton & Kutz, 2023). One particular class of machine learning (ML) approaches called hybrid solvers refers to combining ideas (or full algorithmic blocks) from classical numerical solvers and machine learning (Schenck & Fox, 2018; Um et al., 2020; Kochkov et al., 2021; Jagtap et al., 2022; Lienen & Günnemann, 2022; Karlbauer et al., 2022; Li & Farimani, 2022; Toshev et al., 2023b).
This has been one of the main reasons for the development of differentiable fluid mechanics solvers like PhiFlow (Holl et al., 2020), JAX-CFD (Kochkov et al., 2021), and JAX-Fluids (Bezgin et al., 2023). However, these three frameworks implement Eulerian, i.e., grid-based, solvers, and we lead the way to a JAX-based Lagrangian CFD solver. Eulerian solvers refer to numerical methods that discretize space into static volume elements and then track the evolution of the fluid properties at these positions, while Lagrangian solvers discretize individual material elements, which are then shifted in space following the local velocity field.
@@
As SPH techniques advance, various software packages have arisen in the past few years. However, most of them are designed for high-performance computing (HPC) applications and are typically implemented in low-level languages like C++ or Fortran (Crespo et al., 2015; Koschier et al., 2019; Zhang et al., 2021), with two notable exceptions: PySPH (Ramachandran et al., 2021) in Python and juSPH (Luo et al., 2022) in Julia. Nevertheless, hardly an SPH implementation exists which is out-of-the-box compatible with modern deep learning frameworks like TensorFlow (Abadi et al., 2015), PyTorch [^1] For similar work in progress, we refer the reader to TorchSPH https://github.com/wi-re/pytorchSPH. Compared to our JAX-SPH, this code uses PyTorch, different SPH algorithms, different experiments, and does not validate the solver or the gradients. (Paszke et al., 2019) or JAX (Bradbury et al., 2018), i.e., leverages automatic differentiation to power differentiable solvers like PhiFlow (Holl et al., 2020). Upon identifying the lack of an ML-ready SPH solver, we choose JAX for the framework implementation for two reasons: (a) JAX tends to be faster for operations on graphs even after PyTorch 2.0 has been introduced (see Appendix F in Toshev et al. (2024b)), and (b) we can use the cell list-based neighbor search routine of the JAX-MD library (Schoenholz & Cubuk, 2020). We note that with our solver, we target the easier integration of SPH with ML workflows rather than developing a parallel HPC code, and we exclusively use the AD routine grad by JAX. We leave the implementation of better custom adjoints along the lines of Ma et al. (2021); Kidger (2022); Nadarajah & Jameson (2000) to future work.
In this work, we significantly extend the codebase used for dataset generation within LagrangeBench [^2] https://github.com/tumaer/lagrangebench/blob/main/notebooks/data_gen.ipynb. (Toshev et al., 2024b) and demonstrate the utility of the obtained gradients in multiple ways. Our contributions are:
@@
  • Validating the accuracy of the automatic differentiation-based gradients (Griewank & Walther, 2008) over 5 solver steps.
  • An open-source Python package at pypi.org/project/jax-sph.
@@ 2 SPH Solver
In this section, we introduce the components included in our solver code. The validation of these can be found in Appendix A, for which we include various solver validation cases.
@@
The preferred way of estimating the density (Monaghan, 2005) is through density summation $\rho_{i}=\sum_{j}m_{j}W(r_{ij}|h)$, with $m_{j}$ being the mass of the adjacent particles $j$, $h$ the smoothing length of the kernel, and $r_{ij}$ the interparticle distance. However, this approach leads to unphysically low densities at free surfaces, forcing the use of density evolution, i.e., numerically integrating the mass conservation Eq. 1. We also implement a density field reinitialization method  (Zhang et al., 2017a) to mitigate errors arising from density evolution (Colagrossi & Landrini, 2003). In weakly compressible SPH, the pressure is defined as a function of density through an equation of state, and in our code, we choose the formulation by Monaghan (1994) for standard and transport velocity formulation SPH. The equation of state used for the Riemann SPH formulation is the one proposed by Zhang et al. (2017b).
@@
Riemann SPH. We also add the Riemann SPH solver formulation by Zhang et al. (2017b), which is based on the idea of introducing a simple low dissipation limiter to a classical Riemann solver, resulting in decreased numerical dissipation. We construct a one-dimensional linear inter-particle Riemann problem along a unit vector pointing from particle $i$ to particle $j$, which implicitly regularizes both the momentum equation and the mass conservation. We implement the dissipation limiter as proposed by Zhang et al. (2017b).
Wall boundaries. We follow the generalized wall boundary condition approach by Adami et al. (2012) to enforce different boundary conditions for the standard SPH and the transport velocity SPH formulation. This approach, in essence, implements two physical constraints: (a) impermeability of walls and (b) viscous effects at walls. For impermeability, the pressure of the adjacent fluid particles is assigned to the dummy wall particles, which leads to a zero pressure gradient in the wall-normal direction at the wall surface, and, thus prevents the penetration of fluid particles into the wall. Regarding viscosity, there are two cases to distinguish: (b1) no-slip enforces the no-slip boundary condition, i.e., the fluid must have zero velocity directly at the wall surface, and (b2) free-slip, i.e., the fluid must have a zero velocity normal to a wall, but might have any velocity tangentially. These two conditions are implemented by assigning the inverted velocity of the fluid to the wall particles, either fully for (b1) or only in the wall-normal direction for (b2). See Adami et al. (2012) for more details.
For the Riemann SPH wall boundary implementation, we solve a one-sided Riemann problem as proposed by Zhang et al. (2017b). Unlike the generalized wall boundary condition, this method avoids interpolating and then extrapolating the fluid states for the wall boundary particles. Here, these wall particles are assigned the individual Riemann states depending on the current particle-particle interaction. For a more in-depth explanation, see (Zhang et al., 2017b; Yang et al., 2020).
@@ 3 Experiments
Gradient validation. To test the validity of the gradients obtained through our differentiable solver, we compute analytical solver gradients via automatic differentiation and compare those to numerical gradients from finite difference schemes (Griewank & Walther, 2008). In our setup, gradients are accumulated over 5 solver steps, preceded by 10 (forward only) warm-up steps. Epsilon for finite differences is picked as $50.001dx=5e-$ 5, as smaller values lead to instabilities. Fig. 1 shows the scalar gradients of kinetic energy over position changes $\frac{dE_{kin}}{dr}$ when using 2-dimensional Taylor-Green vortex (TGV) (Brachet et al., 1983; 1984) and lid-driven cavity (LDC) (Ghia et al., 1982) as initial states.
@@
Inverse problem. Our first application case is an inverse problem, representing the class of inverse design and flow control problems that are successfully tackled by differentiable solvers (Holl et al., 2020) as well as differentiable learned solver surrogates (Allen et al., 2022). The scenario involves a 2D box containing a water cube, discretized by 36 particles, which undergoes acceleration due to gravity across 100 solver steps. The task is to find the initial coordinates of the cube given its final state; see Fig. 2.
@@ 4 Conclusion
We have developed JAX-SPH, a framework for simulating Lagrangian fluid problems, which can be easily integrated into design/control problems as well as hybrid solver approaches. By building our code on the high-performance library JAX and validating the simulation results of our solver, we offer a fast and reliable SPH solver in Python. With our work, we hope to accelerate the development of more hybrid Lagrangian solvers, e.g.,  Toshev et al. (2024a), and we leave the addition of more SPH algorithms and simulation cases to future work. One particularly exciting future direction is developing foundation models for PDEs that can operate on both Eulerian and Lagrangian data (Alkin et al., 2024), potentially combined with encoded symmetries (Toshev et al., 2023a).
@@ A.1 Taylor Green Vortex 2D
Fig. 4 shows the absolute velocities of each particle at the start and at the end of the Taylor Green vortex (Brachet et al., 1983; 1984) simulation at $Re=100$. Here, the transport velocity SPH formulation is used. In Fig. 5, one can see the decay of the maximum velocity and the kinetic energy over time for the different methods, standard SPH (SPH), transport velocity formulation SPH (SPH + tvf), and Riemann SPH (Riemann).
(a) $t=0$
(b) $t=5$
Figure 4: 2D Taylor Green vortex velocity magnitudes at the start of the simulation (left) and at $t=5$ (right), calculated using transport velocity formulation SPH.
(a) $u_{max}$
(b) $E_{kin}$
Figure 5: 2D Taylor Green vortex SPH method comparison for $u_{max}$ (left) and $E_{kin}$ (right) between standard SPH, transport velocity formulation SPH, and Riemann SPH at $dx=0.02$ and $dx=0.01$.
@@ A.2 Lid-Driven Cavity 2D
The following figures compare the different methods for a 2D lid-driven cavity (Ghia et al., 1982) simulation at $Re=100$. The reference data for the velocity profiles, i.e., black dots for U (velocity in $x$ direction at a vertical cut through the middle of the cavity) and black squares for V (velocity in $y$ direction at a horizontal cut through the middle of the cavity), are from Ghia et al. (1982).
(a) absolute velocities
(b) velocity profiles
Figure 6: Lid-driven cavity with $dx=0.02$ showing absolute particle velocities of the Riemann solver (left) and velocity profiles of each SPH method at the midsection for U and V (right)
(a) absolute velocities
(b) velocity profiles
Figure 7: Lid-driven cavity with $dx=0.01$ showing absolute particle velocities of the Riemann solver (left) and velocity profiles of each SPH method at the midsection for U and V (right)
@@ A.3 Dam Break 2D
The following Fig. 8 shows the nondimensionalized pressure of a Riemann SPH (Zhang et al., 2017b) dam break (Colagrossi & Landrini, 2003) simulation at different time stamps. The fluid flows from the initial state on the left to the right side, interacts with the wall, and reflects a wave backward throughout the domain. A similar validation plot for the standard SPH formulation of our solver is presented in Appendix B of Toshev et al. (2024b).
@@ A.4 Taylor-Green Vortex 3D
Fig. 9 shows the comparison between the different SPH methods on the 3D TGV, similar to Fig. 5 for 2D. Again, the Reynolds number is set to $Re=100$, and the number of particles in the unit cube is $20^{3}$, $32^{3}$, and $50^{3}$, leading to $dx=0.314$, $dx=0.196$, and $dx=0.126$, respectively. The reference solution is obtained using JAX-Fluids Bezgin et al. (2023) with a $128^{3}$ grid.
(a) $dE/dt$
(b) $E_{kin}$
Figure 9: 3D Taylor Green vortex comparison for $dE/dt$ (left) and $E_{kin}$ (right) between standard SPH, transport velocity formulation SPH, and Riemann SPH with $dx=0.314$, $dx=0.196$, and $dx=0.126$
@@ B.1 Implementation details
Given the number of temporal coarsening steps $M$ and the number of SitL calls n_sitl, each call accounts for $\frac{M}{\texttt{n\_sitl}}$ actual timesteps. This results in two distinct time steps ($dt$), necessitating the transformation of input properties to the SitL physical space.
1def sitl_forward(self, r, u_M_norm, dt, M, n_sitl, sigma_uM, sigma_aM):
@@ 8 M: level of temporal coarsening
14 u_phys = u_M_norm * sigma_uM / (dt * M) # to physical space
15 nbrs_GNN = self.nbrs_GNN_update(r) # GNN neighbor list
16 nbrs_SPH = self.nbrs_SPH_update(r) # SPH neighbor list
@@
34 nbrs_GNN = self.nbrs_GNN_update(r_new)
@@
39 # finite difference to get M-step effective quantities
@@ B.2 Training and results
Solver-in-the-Loop was trained with LagrangeBench (Toshev et al., 2024b), with a custom RPF 2D dataset with 20-step temporal coarsening. For SitL, n_sitl=3. Both the corrector model and the reference GNS (Sanchez-Gonzalez et al., 2020) are message-passing networks with 10 layers and 64 latent dimensions. The starting learning rate is set to $31e-$ 3 for both, and noise_std is set to $51e-$ 5 for SitL and to $31e-$ 3 for GNS. Table 1 shows the LagrangeBench performance metrics on these models. Best models are picked based on the $\text{MSE}_{20}$ loss on the validation set.